# Vortices in the Many-Body Excited States of Interacting Bosons in Two Dimension

### v2

@@ I Introduction
In the present work, we focus on a gas of atoms with short-range interactions inside a 2D circular plaquette. Unlike in 1D, there are no analytical solutions for vortices – neither in the many-body nor in the MF approach. In fact, even numerically, it is not straightforward to find such dynamical solutions to the MF, that would posses a single vortex with core outside the center of the trap. Our starting points are the many-body yrast states – states minimizing energy at fixed momentum, which we find using the exact diagonalization method with importance truncation [28]. We show that during consecutive “measurements” of positions of particles initiated in an yrast state, one reveals vortices. They may appear off the plaquette center, and at random positions, but are robust dynamically, as confirmed by evolving it with the mean-field equations. In a sense, our complicated procedure serves approximate solution to the MF approach. Special attention is paid to states with angular momentum $j\hbar$, with integer $j$ not larger than the total number of atoms $N$. We show that if $j=N$, the position of the vortex revealed from the quantum states fluctuates. As a result, the average density profile differs between the MF and many-body approach, even in this case. The differences are attributed to the quantum fluctuations, as confirmed by the Bogolubov analysis, similarly to the case of "graying" of the 1D black soliton [15]. Most of the attention is paid to the case $j<N$, in which a vortex emerges out of the center.
In Sec. II we present our model. Sec. III is devoted to the ground state, for which MF gives very good results, for as few as $6$ atoms. In Sec. IV we show that vortices are not apparent in the densities, but they emerge in a stochastic process mimicking measurements on many-body excited states. We conclude in Sec V.
@@ II Model
Figure 2: Examples of radial parts of the single particle states used in numerics, Eq. (4) for $m_{j}=0$ (left) and $m_{j}=1$ (right). All radial parts for $m_{j}=0$ are the 00th Bessel functions, but rescaled by $\alpha_{j}$ (with $\alpha_{j}$ being zeros of the 00th bessel function). Similarly, all radial parts for $m_{j}=1$ equal to rescaled 1st Bessel function.
@@
Solutions can be classified into three families: the “$+$ family”, the “-- family”, and the “0 family”. These names correspond to the normalization of the expression $\langle u_{n}|u_{n}\rangle-\langle v_{n}|v_{n}\rangle$, which equals $+1$, $-1$ and 00, respectively [33]. Having $u_{n}(\bm{r})$ and $v_{n}(\bm{r})$ one is in position to compute quantities to characterize the system, as the ones listed below.
@@
where sum of the coefficients $\lambda_{i}$ is equal to $1$. Regime of weak interactions corresponds to situations where majority of the system lies in a single state, say $\psi_{0}$, with its eigenvalue $\lambda_{0}$ being close to $1$, that is $\lambda_{0}\gg\lambda_{i}$ for $i\neq 0$.
@@ III Ground state
Figure 3: Energy per atom for states with $N=6$ particles as a function of coupling constant $g$ obtained from the exact diagonalization (solid lines), GPE approach (black dashed lines) and Bogoliubov approach (red dot-dashed lines). The lower curves correspond to the ground state, i.e. 00 total angular momentum, whereas the upper ones to the states minimizing energy with total angular momentum equal to the total number of particles. All results are presented for $\sigma=0.1\ R$.
@@ IV Yrast states
$G_{N}=\langle\hat{\psi}^{\dagger}(\bm{r}_{1})\ldots\hat{\psi}^{\dagger}(\bar{\bm{r}}_{N-1})\hat{\psi}^{\dagger}(\bm{r})\hat{\psi}(\bm{r})\hat{\psi}(\bar{\bm{r}}_{N-1})\ldots\hat{\psi}(\bar{\bm{r}}_{1})\rangle.$ (20)
@@ V Conclusions
We studied a 2D gas in these excited states that minimize energy at a given angular momentum, known as the yrast states. In a 1D version of this problem, such states are related to mean-field solitons. We have shown that in the 2D version, the states are related to vortices. The vortices emerge spontaneously during measurements, appearing in the high-order correlation functions of the yrast states. When the total angular momentum is less that the number of particles in the system, the system exhibits a single off-center vortex rotating around the center. We demonstrate that the MF-like vortex arising in this manner is always subject to center-of-mass fluctuations, as confirmed by Bogoliubov analysis.
@@ VI Acknowledgements
K.P. acknowledge support from the (Polish) National Science Center Grant No. 2019/34/E/ST2/00289. M.Ś. acknowledge support from the Project no. 2022/45/P/ST3/04237 co-funded by the National Science Centre, Poland and the EU’s H2020 research and innovation programme under the MSCA GA no. 94533.
@@ Appendix B Numerical methods
In this paper we represent the interatomic potential with a narrow Gaussian of range $\sigma=0.1R$. We study up to $6$ atoms – the typical distance between them is ca. $l=\sqrt{\pi R^{2}/6}\approx 0.72R$. As $\sigma$ is substantially smaller than $l$, the interaction potential is practically a short-range one. However, we avoid using the Dirac delta potential directly, as it requires proper regularization in 2D to ensure a finite ground-state energy (see [41] and [42]).

### v3

@@ I Introduction
In the present work, we focus on a gas of atoms with short-range interactions inside a 2D circular plaquette (see Fig. 1). Unlike in 1D, there are no analytical solutions for vortices – neither in the many-body nor in the MF approach. In fact, even numerically, it is not straightforward to find such dynamical solutions to the MF, that would posses a single vortex with core outside the center of the trap. Our starting points are the many-body yrast states – states minimizing energy at fixed momentum, which we find using the exact diagonalization method with importance truncation [28]. We show that during consecutive “measurements” of positions of particles initiated in an yrast state, one reveals vortices. They may appear off the plaquette center, and at random positions, but are robust dynamically, as confirmed by evolving it with the mean-field equations. In a sense, our complicated procedure serves approximate solution to the MF approach. Special attention is paid to states with angular momentum $j\hbar$, with integer $j$ not larger than the total number of atoms $N$. We show that if $j=N$, the position of the vortex revealed from the quantum states fluctuates. As a result, the average density profile differs between the MF and many-body approach, even in this case. The differences are attributed to the quantum fluctuations, as confirmed by the Bogolubov analysis, similarly to the case of "graying" of the 1D black soliton [15]. Most of the attention is paid to the case $j<N$, in which a vortex emerges out of the center.
In Sec. II we present our model. Sec. III is devoted to the ground state, for which MF gives very good results, for as few as 666 atoms. In Sec. IV we show that vortices are not apparent in the densities, but they emerge in a stochastic process mimicking measurements on many-body excited states. We conclude in Sec V.
@@ II Model
Figure 2: Examples of radial parts of the single particle states used in numerics, Eq. (4) for $m_{j}=0$ (left) and $m_{j}=1$ (right). All radial parts for $m_{j}=0$ are the 0th Bessel functions, but rescaled by $\alpha_{j}$ (with $\alpha_{j}$ being zeros of the 0th bessel function). Similarly, all radial parts for $m_{j}=1$ equal to rescaled 1st Bessel function.
@@
Solutions can be classified into three families: the “$+$ family”, the “-- family”, and the “0 family”. These names correspond to the normalization of the expression $\langle u_{n}|u_{n}\rangle-\langle v_{n}|v_{n}\rangle$, which equals +1+1+ 1, 1-1- 1 and 0, respectively [33]. Having $u_{n}(\bm{r})$ and $v_{n}(\bm{r})$ one is in position to compute quantities to characterize the system, as the ones listed below.
@@
where sum of the coefficients $\lambda_{i}$ is equal to 111. Regime of weak interactions corresponds to situations where majority of the system lies in a single state, say $\psi_{0}$, with its eigenvalue $\lambda_{0}$ being close to 111, that is $\lambda_{0}\gg\lambda_{i}$ for $i\neq 0$.
@@ III Ground state
Figure 3: Energy per atom for states with $N=6$ particles as a function of coupling constant $g$ obtained from the exact diagonalization (solid lines), GPE approach (black dashed lines) and Bogoliubov approach (red dot-dashed lines). The lower curves correspond to the ground state, i.e. 0 total angular momentum, whereas the upper ones to the states minimizing energy with total angular momentum equal to the total number of particles. All results are presented for $\sigma=0.1\ R$.
@@ IV Yrast states
$G_{N}=\langle\hat{\psi}^{\dagger}(\bar{\bm{r}}_{1})\ldots\hat{\psi}^{\dagger}(\bar{\bm{r}}_{N-1})\hat{\psi}^{\dagger}(\bm{r})\hat{\psi}(\bm{r})\hat{\psi}(\bar{\bm{r}}_{N-1})\ldots\hat{\psi}(\bar{\bm{r}}_{1})\rangle.$ (20)
@@ V Conclusions
We studied a 2D gas in these excited states that minimize energy at a given angular momentum, known as the yrast states. In a 1D version of this problem, such states are related to mean-field solitons. We have shown that in the 2D version, the states are related to vortices. The vortices emerge spontaneously during measurements, appearing in the high-order correlation functions of the yrast states. When the total angular momentum is less that the number of particles in the system, the system exhibits a single off-center vortex rotating around the center. We demonstrate that the MF-like vortex arising in this manner is always subject to center-of-mass fluctuations, as confirmed by Bogoliubov analysis. The data that support the findings of this article are openly available [41].
@@ VI Acknowledgements
K.P. acknowledge support from the (Polish) National Science Center Grant No. 2019/34/E/ST2/00289. M.Ś. acknowledge support from the Project no. 2022/45/P/ST3/04237 co-funded by the National Science Centre, Poland and the EU’s H2020 research and innovation programme under the MSCA GA no. 945339.
@@ Appendix B Numerical methods
In this paper we represent the interatomic potential with a narrow Gaussian of range $\sigma=0.1R$. We study up to 666 atoms – the typical distance between them is ca. $l=\sqrt{\pi R^{2}/6}\approx 0.72R$. As $\sigma$ is substantially smaller than $l$, the interaction potential is practically a short-range one. However, we avoid using the Dirac delta potential directly, as it requires proper regularization in 2D to ensure a finite ground-state energy (see [42] and [43]).